OpenStudy (diamond77108):

plz help will medal −3−3/4(c−4)=5/4 answers 2 -5/3 -15/16 -11

OpenStudy (kittiwitti1):

\[\frac{-3-3}{4(c-4)}=\frac{5}{4}\]first let's try to get the part with \(c\) alone, so:\[(-3-3)\div\frac{5}{4}=4(c-4)\]

OpenStudy (kittiwitti1):

-3-3 is -6:\[-6\div\frac{5}{4}=4(c-4)\]Dividing by a fraction also means multiplying by its reciprocal (the fraction flipped around):\[-6\times\frac{4}{5}=4(c-4)\]\[\frac{-24}{5}=4(c-4)\]

OpenStudy (kittiwitti1):

Distributing the 4(c-4):\[-\frac{24}{5}=4c-4(4)\]\[-\frac{24}{5}=4c-16\]\[-\frac{24}{5}+16=4c\]

OpenStudy (kittiwitti1):

Now isolate the \(c\). A hint... \(16=\large{\frac{80}{5}}\)

OpenStudy (diamond77108):

16

OpenStudy (diamond77108):

16=16

OpenStudy (kittiwitti1):

What?

OpenStudy (kittiwitti1):

You can't add -24/5 and 16 without converting 16 to fractional form.

OpenStudy (diamond77108):

-24/5+4=c i really dont get the way you doing it

OpenStudy (kittiwitti1):

Is this the equation?\[\frac{-3-3}{4(c-4)}=\frac{5}{4}\]

OpenStudy (diamond77108):

\[-3- \frac{ 3 }{ 4}(c-4)= \frac{ 5 }{ 4 }

OpenStudy (kittiwitti1):

I guess you could also do this:\[(-3-3)\times4=5\times[4(c-4)]\]

OpenStudy (kittiwitti1):

You can always do cross multiplication I guess. I'm not sure what your original problem was so I can't really help you.

OpenStudy (diamond77108):

\[-3-\frac{ 3 }{ 4 }(c-4)=\frac{ 5 }{ 4}\] is my problem

OpenStudy (mathstudent55):

My suggestion is to first multiply both sides by 4 to get rid of all fractions.

OpenStudy (mathstudent55):

Can you do that and show what you get?

OpenStudy (diamond77108):

but im stiil gettin fractions

OpenStudy (mathstudent55):

Let me show you.

OpenStudy (mathstudent55):

\(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=\color{red}{4}(\dfrac{ 5 }{ 4})\) See the red 4 on each side? Each side is being multiplied by 4.

OpenStudy (mathstudent55):

The right side is simply 4 * 5/4, which is 4/1 * 5/4. \(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=\color{red}{\dfrac{4}{1}}(\dfrac{ 5 }{ 4})\) \(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=\color{red}{\dfrac{\cancel 4}{1}}(\dfrac{ 5 }{ \cancel 4})\) \(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=\color{red} 5\) You see how the right side simplifies to only 5?

OpenStudy (diamond77108):

so dose 3/4 become 3 to

OpenStudy (mathstudent55):

Yes, 3/4 becomes just 3 when you multiply by 4, but on the left side, you are multiplying 4 by a quantity, so you need to use the distributive property.

OpenStudy (mathstudent55):

Let's work on the left side now. \(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=5 \)

OpenStudy (diamond77108):

so the equation is -3-3(c-4)=5

OpenStudy (mathstudent55):

\(\color{red}{4}[-3-\dfrac{ 3 }{ 4 }(c-4)]=5 \) Distribute the 4: \(\color{red}{4}(-3)-\color{red} 4 \times\dfrac{ 3 }{ 4 }(c-4)]=5 \) See how the 4 multiplies both terms?

OpenStudy (mathstudent55):

Now we do each multiplication by 4. You did the second one correctly, but you can;t forget the first one.

OpenStudy (diamond77108):

-12-3(c-4)=5 this

OpenStudy (mathstudent55):

\(\color{red}{4}(-3)-\color{red} 4 \times\dfrac{ 3 }{ 4 }(c-4)=5 \) \(\color{red}{-12}-\color{red}{3}(c-4)=5 \)

OpenStudy (mathstudent55):

Correct. Now you need to distribute the -3 through the parentheses.

OpenStudy (diamond77108):

\[-12-3c-12=5\]

OpenStudy (mathstudent55):

Be careful with the signs. You are multiplying each number in the parentheses by NEGATIVE 3.

OpenStudy (mathstudent55):

\(-12 - 3(c - 4) = 5\) \(-12 - 3[c + (- 4)] = 5\) \(-12 - 3 \times c + (-3)(-4) = 5\)

OpenStudy (mathstudent55):

\(-12 - 3 c + 12 = 5\)

OpenStudy (mathstudent55):

-3 * c = -3c -3 * (-4) = +12 <--- be careful with this multiplication

OpenStudy (diamond77108):

ok yeah cuz negative * negative = positive

OpenStudy (mathstudent55):

correct

OpenStudy (mathstudent55):

Now, combine like terms on the left side. Add -12 and 12 on the left side. What do you get?

OpenStudy (diamond77108):

0

OpenStudy (diamond77108):

they cancel out

OpenStudy (mathstudent55):

Good. That is correct. You have this left: \(- 3 c = 5\) Solve for c.

OpenStudy (diamond77108):

c=-5/3

OpenStudy (mathstudent55):

Correct.

OpenStudy (diamond77108):

thx

OpenStudy (mathstudent55):

yw